Choose the graph for the upper 10% for a standard normal distribution FON O L L トラッ FON NO
Answer #1

Here Z is the standard normal random variable

By looking at the picture it is likely that 3rd option is our answer so, we check by calculating probability for the 3rd option

Probability of Z greater than 1.28 = P(Z>1.28)

P(Z>1.28) = 1 - P(Z<1.28)

P(Z>1.28) = 1 - 0.8997 (we get P(Z<1.28)= 0.8997 from z-table)

P(Z>1.28) = 0.1003

So, Percentage of data = 0.1003 \small \times 100% = 10.03% = 10% (rounded)

Now we can confirm that option 3 is our answer as option 3 represent upper 10% of the standard normal distribution

Answer : 3rd Option

O -1 FON 128 1